OpenStudy (anonymous):

write the expression as a single lograrithm, Express powers as factors. ln((x/x-7))+ln((x+7/x))-ln(x^2-49)=______________

OpenStudy (anonymous):

\[\ln \frac{ a }{ b }=\ln a - \ln b\]\[\ln ab = \ln a + \ln b\]Factor x^2 -49 and you can do it easily.

OpenStudy (anonymous):

would this be it \[(\ln(x)-\ln(x-7))+(\ln(x+7)-\ln(x))-(\ln(x+7)(x-7))\] @Anonymous1921

OpenStudy (anonymous):

Yes. And ln[(x+7)(x-7)]=ln(x+7)+ln(x-7). So you can cancel out something.

OpenStudy (anonymous):

okay so it will look like wouldn't everything be cancel?

OpenStudy (anonymous):

@Anonymous1921

OpenStudy (anonymous):

I think -2ln(x-7) would be left. Check your calculation again. I might be wrong.

OpenStudy (anonymous):

so I got it to look like lnx-lnx-7+lnx+7-lnx-lnx-7+lnx+7

OpenStudy (anonymous):

then the lnx would be cancel and the lnx+7 would be cancel leaving -2lnx-7 ?

OpenStudy (anonymous):

ln(x)-ln(x-7)+ln(x+7)-ln(x)-ln(x+7)-ln(x-7). Oh, sorry. Everything would be cancelled out.

OpenStudy (anonymous):

really?? but the ln(x-7) doesnt get cancel out

OpenStudy (anonymous):

so wouldnt it be -2ln(x-7)

OpenStudy (anonymous):

Oh, yeah. -2ln(x-7).

OpenStudy (anonymous):

okay thanks you for the big help!

OpenStudy (anonymous):

You're welcome :)